OpenStudy (anonymous):

How many moles of carbon dioxide gas can be produced when 12.8 moles of benzene (C6H6) react with excess oxygen? Unbalanced equation: C6H6 + O2 → CO2 + H2O can someone help me with this im so confused on how to do it?

OpenStudy (anonymous):

2 C6H6 + 15 O2 = 12 CO2 + 6 H2O i already know the balanced equation

OpenStudy (matt101):

Balanced reaction looks good. To figure out the answer to this question, we need to set up a ratio and solve. We want to find the amount of CO2 produced given the amount of C6H6 reacted. According to the balanced reaction, for every 2 moles of C6H6 reacted, 12 moles of CO2 will be produced. If 4 moles of C6H6 reacted, you'd get 24 moles of CO2 produced. If 8 moles of C6H6 reacted, you'd get 48 moles of CO2 produced. They always react in the same ratio, as you can see: 2 C6H6:12 CO2. We're given a funny amount of C6H6 to start, but we can find the answer easily using this ratio: \[\frac {2 \space C_6H_6}{12 \space CO_2}=\frac{12.8 \space C_6H_6}{x \space CO_2}\] Since we know these guys always react in the same ratio, just solve this equation for x to find the moles of CO2 produced!

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

actually wait...where did the 12.8 c6h6 / x co2 come from?

OpenStudy (anonymous):

@matt101

OpenStudy (matt101):

It's just matching the ratio we're given in the reaction. Basically the equation says: If 2 moles of C6H6 react, 12 moles of CO2 are produced. That means if 12.8 moles of C6H6 react, ____ moles of CO2 are produced. You're solving for the blank. If it's any easier, you can get rid of the words and just use the numbers in the equation. Solve for x and post your steps here!

OpenStudy (anonymous):

81.6?

OpenStudy (anonymous):

12 x 6.8=81.6 2=12 4=24 6=36 8=48 10=60 12= 72 12.8=81?

OpenStudy (anonymous):

am i right or wrong, i kinda feel like im wrong.

OpenStudy (matt101):

All you're doing is solving for x! First cross-multiply what's in the equation I wrote above so you get: 2x=12.8(12) 2x=153.6 Next just divide both sides by 2 to isolate x: x=76.8 Since x=76.8, if you react 12.8 moles of C6H6, you will produce 76.8 moles of CO2! Does that make sense?

OpenStudy (anonymous):

ohhhhh ok that makes sense, thank you!